OpenStudy (anonymous):

help me with this question given that g(y) = y^2-1, Find g(a-1).

OpenStudy (anonymous):

=(a-1)^2 -1 =a(a-2)

OpenStudy (anonymous):

put (a-1) in place of y (a-1)^2-1=a^2-2a+1-1=a^2-2a

OpenStudy (anonymous):

thanks alot..:0

hero (hero):

At least Arnab09 provided some kind of explanation. I would have rather given the poster a chance to solve it himself. Saying " put (a-1) in place of y" was a good enough hint.

OpenStudy (anonymous):

G(x)=5x-2 Sub y for f(x) y=5x-2 Switch the x and y x=5y-2 Solve for y x+2=5y 1/5x+2/5=y Sub f^-1(x) for y f^-1(x)=1/5x+2/5

hero (hero):

@Rohangrr , Is there a meaningful reason for post that?

OpenStudy (anonymous):

@Hero yes